What is the total time of flight of an object projected vertically upwards with a speed of 30ms? (g 10 ms'2)
1
Expert's answer
2021-08-19T14:21:46-0400

Gives

v=30m/sec

g=10m/sec2

Total time period of flight

"T=\frac{2usin\theta}{g}"

"T_{max}=\frac{2u}{g}\\sin\theta=sin90°=1"

Put value

"T=\frac{2u}{g}=\frac{2\times30}{10}=6sec"